OpenStudy (anonymous):

Mike needs $41,300 to purchase a boat. How much money will he need to invest today in a savings account earning 3.7% interest, compounding quarterly, to have enough money to purchase the boat in 15 years?

OpenStudy (anonymous):

$4,669.01 $39,071.30 $35,624.90 $23,769.75

hartnn (hartnn):

the formula for compound interest is : \(I = P (1+r/4)^{4n} -P\) where P is the principal amount, n= time period in years and r= rate of interest.

OpenStudy (anonymous):

p= 41,300 r= 0.00925 right?

hartnn (hartnn):

and the final amount is say, A = I+P = \((1+r/4)^{4n} \) not really, he needs final amount as 41300 , so A = 41,300

hartnn (hartnn):

r= 3.7% = 0.037

OpenStudy (anonymous):

but you divide r/4 right?

OpenStudy (anonymous):

so whats P?

hartnn (hartnn):

so, r/4 = 0.037/4=0.00925

hartnn (hartnn):

P is money invested today which you need to find out

OpenStudy (anonymous):

oh ok. so how do i make that a formula?

hartnn (hartnn):

you plug in all the values first, n= 15

OpenStudy (anonymous):

its nt right? so 15x4=60

hartnn (hartnn):

the ecponent = 4n = 60 yes

OpenStudy (anonymous):

ok so whats next?

hartnn (hartnn):

what u got after plugging in ?

OpenStudy (anonymous):

41300(1+0.00925)^60?

hartnn (hartnn):

41300 = P (1+0.00925)^60

hartnn (hartnn):

use calculator for this (1+0.00925)^60

OpenStudy (anonymous):

1.7375

hartnn (hartnn):

yes, 41300 = P *1.7375 now can you find P ?

OpenStudy (anonymous):

do i divide 41300/1.7375

hartnn (hartnn):

yes

OpenStudy (anonymous):

awesome so i got 23,769.78

OpenStudy (anonymous):

the answer is off by a few cents but that doesnt really matter, right?

hartnn (hartnn):

i too get the same :)

hartnn (hartnn):

yes, it doesn't. i clculated it and got 23769.747

OpenStudy (anonymous):

ok cool. can you help me with something else please?

hartnn (hartnn):

sure, :)

OpenStudy (anonymous):

thanks :) ok so Julianne wants to take a trip around the world. She plans to deposit $175 at the beginning of each month into an investment with a 3.25% interest rate, compounded monthly. How much will she have in the account after 12 years? 30,770.00 $30,853.33 $533,286.11 $550,617.91

hartnn (hartnn):

\(A= P (1+r/2)^{2n} \)

OpenStudy (anonymous):

same as before?

hartnn (hartnn):

the formula changed for monthly...

hartnn (hartnn):

now you have to find A

OpenStudy (anonymous):

P=175?

hartnn (hartnn):

yes, r = 3.25 % = 0.0325

hartnn (hartnn):

n=12

OpenStudy (anonymous):

r=0.002708

hartnn (hartnn):

why ?

OpenStudy (anonymous):

because its divided by 12

hartnn (hartnn):

its r/2 , not r/12

OpenStudy (phi):

If you are making monthly deposits you use a more complicated formula FV = PMT * ( ( (1 + i)^n - 1) / i )

hartnn (hartnn):

oh, wait i am sorry, its r/12 only

hartnn (hartnn):

\(A= P (1+r/12)^{12n}\)

OpenStudy (anonymous):

oh ok. @phi what are the values of PMT?

OpenStudy (phi):

PMT is the payment $175

OpenStudy (anonymous):

@hartnn its 12x12=144 for nt, right?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

ok so 175(1+0.002708)^144?

hartnn (hartnn):

i believe that would give you the answer...

OpenStudy (anonymous):

im getting 258.34

hartnn (hartnn):

me too

OpenStudy (phi):

this problem is different from the 1st one. You are making regular deposits every month of 175 even with no interest you would expect 175*144 with compounded monthly interest, you have to use the formula I posted up above

OpenStudy (anonymous):

so i multiply 175*144 and i get 25200.

OpenStudy (anonymous):

would that be the PMT?

OpenStudy (phi):

\[ FV = pymt\cdot \frac{ (1+i)^n -1}{i} \]

OpenStudy (phi):

where i is the monthly interest 0.0325/12 and n is in months 12years * 12 months/year = 144

OpenStudy (anonymous):

whats the payment?

OpenStudy (phi):

She plans to deposit $175 at the beginning of each month

OpenStudy (anonymous):

so i multiply 175 by 144?

OpenStudy (phi):

no, you have to use the formula. Do it step by step. \[FV = pymt\cdot \frac{ (1+i)^n -1}{i} \] replace all the letters with numbers: you know pymt is 175 you know i and you you know n

OpenStudy (anonymous):

im not getting it

OpenStudy (phi):

First step, what is i for this problem ?

OpenStudy (anonymous):

0.0325/12= 0.002708333

OpenStudy (phi):

so we put that into the formula to get \[ FV = pymt\cdot \frac{ (1+i)^n -1}{i} \] \[ FV = pymt\cdot \frac{ (1+0.002708333)^n -1}{0.002708333} \]

OpenStudy (phi):

what is n? (in months) what is pymt ?

OpenStudy (anonymous):

n=12 pymt= 175

OpenStudy (phi):

remember, the term is 12 years, monthly it is 144 months n is 144 so now we have \[ FV =175\cdot \frac{ (1.002708333)^{144} -1}{0.002708333} \] notice I simplified 1+0.002708333 to 1.002708333

OpenStudy (phi):

now time to get out the calculator

OpenStudy (anonymous):

30,769.99 rounded up is 30,770 :)

OpenStudy (phi):

yes, success!

OpenStudy (anonymous):

lol thanks. do you mind helping some more?

OpenStudy (anonymous):

aww ok. well thanks so much.